Find the points on the cone z2 =x2 + y2 that are closest tothe point (6, 2,0).
Answer #1
assume the required point is Q(x, y, z), or Q(x, y,±(x2 + y2))
the square of the distance between Q and P(6, 2, 0) is
F(x, y) = (x - 6)2 + (y - 2)2 + x2+ y2)
= 2x2 + 2y2 - 12x - 4y + 40
= 2(x2 + y2 - 6x - 2y + 20)
Fx = 2(2x - 6) = 0, x = 3
Fy = 2(2y - 2) = 0, y = 1
so Q is (3, 1, ±√10)